OpenStudy (anonymous):

Can someone please help with this problem? Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1, 3) m = 3

OpenStudy (texaschic101):

y = mx + b slope(m) = 3 (1,3) x = 1 and y = 3 now we sub.... 3 = 3(1) + b 3 = 3 + b 3 - 3 = b 0 = b so we have : y = 3x + 0 or this way... y - y1 = m(x - x1) slope(m) = 3 (1,3) x1 = 1 and y1 = 3 sub y - 3 = 3(x - 1) y - 3 = 3x - 3 y = 3x - 3 + 3 y = 3x + 0

OpenStudy (anonymous):

Thank you so much!!